OpenStudy (anonymous):

lim x->64 √(x) -8)/x-64 I got 0 but the answer is 1/16 how?

OpenStudy (anonymous):

\[\frac{\sqrt{x}-8}{x-64}=\frac{\sqrt{x}-8}{(\sqrt{x}-8)(\sqrt{x}+8)}\] cancel, replace \(x\) by \(64\)

OpenStudy (anonymous):

Smh I didn't see it like that at all thank you!

OpenStudy (anonymous):

be careful, because if you replace \(x\) by 64 you do not get 0, but rather you get \(\frac{0}{0}\) which is undefined. that means you have to do more work, and usually it involves factoring and canceling

OpenStudy (anonymous):

yw